# Online regularization of Poincaré map of storage rings with Shannon entropy

### v1

@@ Abstract
Shannon Entropy is adopted to quantify the chaos of measured Poincaré maps in the National Synchrotron Light Source-II (NSLS-II) storage ring. The recurrent Poincaré maps, constructed from beam position monitor’s turn-by-turn readings, are commonly used to observe the nonlinearity in ring-based accelerators. However, these observations typically only provide a qualitative observation. With some canonical transformations on Poincaré maps, not only can the commonly used nonlinear characterizations be extracted, but more importantly, the chaos can be quantitatively measured with entropy. Entropy, therefore as a chaos indicator, is used for online Poincaré map regularization and dynamic aperture optimization in the NSLS-II ring.
@@
In information theory, Shannon entropy Shannon (1948) is used to quantify the uncertainty or unpredictability in a random variable. Due to the common unpredictable behavior in nonlinear dynamical systems, it can be applied to study the chaos Letellier (2006). In a Poincaré map formed with a particle’s $N$-turn coordinate $\mathbf{X}=(x,p_{x},y,p_{y},z,p_{z})^{T}$, the entropy is defined as
@@
A linear regular motion on an invariant torus is completely predicable, its entropy should be zero naturally. But if computing with Eq. (1) in the conventional position-momentum coordinate, the entropy unnecessarily depends on $N$. To address this limitation, we propose applying the following canonical transformations prior to estimating entropy. Assuming the initial condition at a specific longitudinal location is $\mathbf{X}_{0}$. After a full turn it becomes $\mathbf{X}_{1}=\mathbf{MX}_{0}$, with the one-turn transportation matrix $\mathbf{M}$. If the motion is stable, a set of Twiss parameters can be defined to normalize the coordinates into a set of conjugate action-angle variables $(J_{x},\phi_{x},J_{y},\phi_{y},J_{z},\phi_{z})$ Courant and Snyder (1958). In this action-angle coordinate, the linear motion is purely rotational. It has fixed radii $J_{x,y,z}$ and phase advances $\Delta\phi_{x,y,z}=2\pi\nu_{x,y,z}$ in each plane. $\nu_{x,y,z}$ are known as the linear tunes. Now we introduce a parallel-shift transformation to each angle variable, the new variables read as
@@
where $n=0,1,2,\cdots$ is the index of turns, and $2\pi\overline{\nu}$ is the average phase advance per turn. By applying the above transformations, a linear motion is mapped to a single state, with zero entropy correspondingly. If a linear coupling exists there, a decoupling process through similar transformations such as Edward-Teng’s parameterization Edwards and Teng (1973), would be needed first.
When nonlinear magnets are added to a ring lattice, the particle’s coordinate typically do not precisely overlap its previous state due to nonlinear effects such as TSWA and action-angle smearing. The TSWA $(\frac{\partial\nu}{\partial J})$ describes an average phase-advance shift with action amplitude. The distortions of action-angles are quantified by the smear parameters Chao et al. (1988), i.e., their fractional deviations $(\frac{\Delta J}{J},\,\frac{\Delta\phi}{\phi})$. From a given initial condition, a new state is evolved after each turn, with the exception of some special points such as the origin and fixed points. Its Poincaré map gradually diffuses with the number of passes. According to the Kolmogorov-Arnold-Moser theorem, when the fractional tune is irrational, the quasi-periodical motion of the particle is distorted but not destroyed. In such instances, a normal form technique Bazzani et al. (1994) is used to transfer the map into a particular simple form. Normal form is a generalized Courant-Synder normalization Courant and Snyder (1958) through a nonlinear canonical transformation. The new map usually has explicit invariants and more symmetries. Therefore, for nonlinear motions, prior to the parallel-shift transformation of Eq. (2), we also implement another normal form with a type-2 generating function
@@
here $i$ is the index of kick amplitudes, $S_{0}$ represents the entropy of the initial lattice configuration used as the reference. The reference is used here because we would like to see the relative entropy change. Once the entropy converges within its measurement precision, the optimization is terminated. The optimal configuration is checked by re-measuring the Poincaré maps and comparing against the initial settings. With the optimal sextupole configuration, the entropy at different amplitudes reduces simultaneously as illustrated in Fig. 3. A significant improvement on the regularity of the maps is achieved by a nonlinear optimization in the vicinity of $\nu_{x}=\frac{2}{9}$ resonance, as shown in Fig. 4. In principle, the reduction of entropy is equivalent to a regularization on the Poincaré map. Consequently upon this regularization, the off-axis injection efficiency is improved from 90% to 100%.
@@
Acknowledgments: This research is supported by the U.S. Department of Energy under Contract No. DE-SC0012704. Coauthor KA is partially sponsored by Brookhaven National Laboratory Graduate Research Internship Program 2024.

### v2

@@ Abstract
Shannon entropy, as a chaos indicator, is used for online Poincaré map regularization and dynamic aperture optimization in the National Synchrotron Light Source-II (NSLS-II) ring. Although various chaos indicators are widely used in studying nonlinear dynamical systems, including modern particle accelerators, it is the first time to use a measurable one in a real-world machine for online nonlinear optimization. Poincaré maps, constructed with the turn-by-turn beam trajectory readings from beam position monitors, are commonly used to observe the nonlinearity in ring-based accelerators. However, such observations typically only provide a qualitative interpretation. We analyze their entropy to quantify the chaos in measured Poincaré maps. After some canonical transformations on the Poincaré maps, not only can the commonly used nonlinear characterizations be extracted, but more importantly, the chaos can be quantitatively calibrated with Shannon entropy, and then used as the online optimization objectives.
@@
Methods
In information theory, Shannon entropy Shannon (1948) is used to quantify the uncertainty or unpredictability in a random variable. Due to the common unpredictable behavior in nonlinear dynamical systems, entropy can be applied to study the chaos Letellier (2006). In a Poincaré map formed with a particle’s $N$-turn coordinate $\mathbf{X}=(x,p_{x},y,p_{y},z,p_{z})^{T}$, the entropy is defined as
@@ Methods
Since a linear regular motion on an invariant torus is completely predicable, its entropy should be zero naturally. But if computing with Eq. (1) in the conventional position-momentum coordinate, the entropy unnecessarily depends on $N$. To address this limitation, we propose applying the following canonical transformations prior to estimating entropy. Assuming the initial condition at a specific longitudinal location is $\mathbf{X}_{0}$. After a full turn it becomes $\mathbf{X}_{1}=\mathbf{MX}_{0}$, with the one-turn transportation matrix $\mathbf{M}$. If the motion is stable, a set of Twiss parameters can be defined to normalize the coordinates into a set of conjugate action-angle variables $(J_{x},\phi_{x},J_{y},\phi_{y},J_{z},\phi_{z})$ Courant and Snyder (1958). In this action-angle coordinate, the linear motion is purely rotational. It has fixed radii $J_{x,y,z}$ and phase advances $\Delta\phi_{x,y,z}=2\pi\nu_{x,y,z}$ in each plane. $\nu_{x,y,z}$ are known as the linear tunes. Now we introduce a parallel-shift transformation to each angle variable, the new variables read as
@@
where $n=0,1,2,\cdots$ is the index of turns, and $2\pi\overline{\nu}$ is the average phase advance per turn. By applying the above transformations, a linear motion is mapped to a single state, with zero entropy correspondingly. If a linear coupling exists there, a decoupling process through similar transformations such as Edward-Teng’s parameterization Edwards and Teng (1973), would be needed first. In principle, the transverse Poincaré maps in accelerators are 4D tori. Since the coupling is weak, or can be decoupled in our study, we can use their 2D projections in the horizontal and vertical planes respectively.
When nonlinear magnets are added to a ring lattice, the particle’s coordinate typically does not precisely overlap its previous state due to nonlinear effects such as TSWA and action-angle smearing. The TSWA $(\frac{\partial\nu}{\partial J})$ describes an average phase-advance shift with action amplitude. The distortions of action-angles are quantified by the smear parameters Chao et al. (1988), i.e., their fractional deviations $(\frac{\Delta J}{J},\,\frac{\Delta\phi}{\phi})$. From a given initial condition, a new state is evolved after each turn, with the exception of some special points such as the origin and fixed points. Its Poincaré map gradually diffuses with the number of passes. According to the Kolmogorov-Arnold-Moser theorem, when the fractional tune is irrational, the quasi-periodic motion of the particle is distorted but not destroyed. In such instances, a normal form technique Bazzani et al. (1994) is used to transfer the map into a particular simple form. Normal form is a generalized Courant-Synder normalization Courant and Snyder (1958) through a nonlinear canonical transformation. The new map usually has explicit invariants and more symmetries. Therefore, for nonlinear motions, prior to the parallel-shift transformation of Eq. (2), we also implement another normal form with a type-2 generating function
@@ Results
here $i$ is the index of kick amplitudes, $S_{0}$ represents the entropy of the initial lattice configuration used as the reference. The reference is used here because we would like to see the relative entropy change. Once the entropy converges within its measurement precision, the optimization is terminated. The optimal configuration is checked by re-measuring the Poincaré maps and comparing against the initial settings. With the optimal sextupole configuration, the entropy at different amplitudes reduces simultaneously as illustrated in Fig. 3. A significant improvement on the regularity of the maps is achieved as shown in Fig. 4. Particularly the irregular motion in the vicinity of $\nu_{x}=\frac{2}{9}$ resonance is mitigated. In principle, the reduction of entropy is equivalent to a regularization on the Poincaré map. Consequently upon this regularization, the off-axis injection efficiency is improved from 90% to 100%.
@@
Most existing ring-based accelerator lattices (including NSLS-II) are non-integrable, therefore they only have approximate invariants when particle oscillation amplitude is sufficiently small. If a ring lattice is integrable and has analytical invariants Danilov and Nagaitsev (2010), the near-zero Shannon entropy might be found through some canonical transformations and be potentially used as the experimental criteria of its integrability. Some further exploration on applying Shannon entropy analysis to nonlinear dynamics might be interesting.
@@
Acknowledgments
We would like to thank Sergei Nagaitsev and Ivan Morozov for their in-depth discussion. This research is supported by the U.S. Department of Energy under Contract No. DE-SC0012704, the NSLS-II Facility Improvement Projects and the DOE HEP award DE-SC0023722. Coauthor KA is partially sponsored by Brookhaven National Laboratory Graduate Research Internship Program 2024.
Author contributions
YL proposed and developed the method, performed the experimental study and prepared the manuscript. DX and YH developed the method and prepared the manuscript. KA and MS performed the experimental study. KH and YH supported the experimental study. RR prepared the manuscript. VS and TS sponsored KA’s internship.
Competing interests
The authors declare no competing interests.